# HarmonyCell: Automating Single-Cell Perturbation Modeling under Semantic and Distribution Shifts

### v1

@@ 1 Introduction
Existing approaches still fall short of the end-to-end goal: turning a user’s scientific intent into a reusable workflow that adapts to heterogeneous single-cell data. As shown in Fig. 1, search-centric deep research agents (Hu et al., 2025; Chai et al., 2025) can retrieve relevant papers and heuristics, but rarely compile them into robust, executable pipelines that ingest messy single-cell data and reliably produce models. General code-centric AutoML agents (Du et al., 2025; Jiang et al., 2025; Ou et al., 2025) often lack biological priors and degrade into reactive trial-and-error, treating each request as a one-off coding task rather than a structured research process. Closest to our setting, CellForge (Tang et al., 2025) shows that multi-agents can design and implement virtual-cell models from raw data and task descriptions; however, it assumes an over-simplified setting and fails to address the dual heterogeneity challenge. Specifically, it lacks mechanisms to (i) automatically harmonize divergent metadata schemas into a canonical interface, and (ii) systematically optimize structural inductive biases to handle complex distribution shifts (e.g., OOD tasks). Consequently, researchers still lack an agent that can autonomously bridge the gap between raw, heterogeneous data and robust model deployment in the fragmented single-cell ecosystem.
@@ 2.2 Coding and Automated Machine Learning Agents
The intersection of Large Language Models (LLMs) and software engineering has facilitated the development of autonomous agents capable of executing end-to-end machine learning workflows (Trirat et al., 2024; Wang et al., 2024). Building upon traditional Automated Machine Learning (AutoML) techniques, recent multi-agent systems leverage the search (Hu et al., 2025; Chai et al., 2025) and code-generation (Zhuge et al., 2024; Wang et al., 2025; Yang et al., 2025; Ou et al., 2025) capabilities of LLMs to automate pipeline construction, ranging from data preprocessing and feature engineering to model architecture search. Recently, some multi-agent systems enables tree-search methods to explore possible solution spaces (Jiang et al., 2025; Du et al., 2025), which accelerate the discovery of scientific problems and coding solutions.
@@ 2.3 AI Scientist: Agents in Scientific Discovery
Beyond code execution, a distinct class of agents, often termed ”AI Scientists,” has been developed recently to automate high-level cognitive tasks inherent to the scientific process both in general-purpose domain (Chai et al., 2025; Yamada et al., 2025; Lu et al., 2024; Ruan et al., 2024) and specific fields (Ruan et al., 2024; Jin et al., 2025; Ni et al., 2024; Zhang et al., 2025; Gottweis et al., 2025; Swanson et al., 2024; Xiao et al., 2024). These AI scientists integrate literature retrieval, knowledge extraction, hypothesis generation and code execution to assist researchers in navigating vast scientific corpora and formulating theoretical insights. In agentic virtual cell modeling, CellForge (Tang et al., 2025) first merge search and code-generation capabilities of LLMs to automate the design of virtual cell models.

### v2

@@ 1 Introduction
Existing approaches still fall short of the end-to-end goal: turning a user’s scientific intent into a reusable workflow that adapts to heterogeneous single-cell data. As shown in Fig. 1, search-centric deep research agents (Hu et al., 2025b; Chai et al., 2025) can retrieve relevant papers and heuristics, but rarely compile them into robust, executable pipelines that ingest messy single-cell data and reliably produce models. General code-centric AutoML agents (Du et al., 2025; Jiang et al., 2025; Ou et al., 2025) often lack biological priors and degrade into reactive trial-and-error, treating each request as a one-off coding task rather than a structured research process. Closest to our setting, CellForge (Tang et al., 2025) shows that multi-agents can design and implement virtual-cell models from raw data and task descriptions; however, it assumes an over-simplified setting and fails to address the dual heterogeneity challenge. Specifically, it lacks mechanisms to (i) automatically harmonize divergent metadata schemas into a canonical interface, and (ii) systematically optimize structural inductive biases to handle complex distribution shifts (e.g., OOD tasks). Consequently, researchers still lack an agent that can autonomously bridge the gap between raw, heterogeneous data and robust model deployment in the fragmented single-cell ecosystem.
@@ 2.2 Coding and Automated Machine Learning Agents
The intersection of Large Language Models (LLMs) and software engineering has facilitated the development of autonomous agents capable of executing end-to-end machine learning workflows (Trirat et al., 2024; Wang et al., 2024). Building upon traditional Automated Machine Learning (AutoML) techniques, recent multi-agent systems leverage the search (Hu et al., 2025b; Chai et al., 2025) and code-generation (Zhuge et al., 2024; Wang et al., 2025; Yang et al., 2025; Ou et al., 2025) capabilities of LLMs to automate pipeline construction, ranging from data preprocessing and feature engineering to model architecture search. Recently, some multi-agent systems enables tree-search methods to explore possible solution spaces (Jiang et al., 2025; Du et al., 2025), which accelerate the discovery of scientific problems and coding solutions.
@@ 2.3 AI Scientist: Agents in Scientific Discovery
Beyond code execution, a distinct class of agents, often termed ”AI Scientists,” has been developed recently to automate high-level cognitive tasks inherent to the scientific process both in general-purpose domain (Wei et al., 2025a; Chai et al., 2025; Yamada et al., 2025; Lu et al., 2024; Ruan et al., 2024; Xu et al., 2025) and specific fields (Ruan et al., 2024; Jin et al., 2025; Ni et al., 2024; Zhang et al., 2025; Gottweis et al., 2025; Swanson et al., 2024; Xiao et al., 2024). By using various methods like tree-structure searching(Wei et al., 2025b), RAG methods(Wei et al., ) and data contruction (Hu et al., 2025a), these AI scientists integrate literature retrieval, knowledge extraction, hypothesis generation and code execution to assist researchers in navigating vast scientific corpora and formulating theoretical insights. In agentic virtual cell modeling, CellForge (Tang et al., 2025) first merge search and code-generation capabilities of LLMs to automate the design of virtual cell models.